# Identity-Preserving Image-to-Video Generation via Reward-Guided Optimization

### v1

@@ 1 Introduction
Identity preservation is a widely popular and explored topic in text-to-video (T2V) generation. Numerous recent methods (He et al., 2024; Liu et al., 2025c; Yuan et al., 2025; Wang et al., 2025; Li et al., 2024; 2025; Song et al., 2025; Xie et al., 2025; Zhang et al., 2025; Zhong et al., 2025; Jiang et al., 2025) attempt to generate high-fidelity videos that maintain consistent identity from user-provided reference subjects, often by incorporating auxiliary identity modules and carefully curated datasets. However, it remains a persistent yet underexplored challenge in image-to-video (I2V) generation. Despite diffusion transformer-based models such as CogVideoX (Yang et al., 2024), HunyuanVideo (Kong et al., 2024), and Wan (Wan et al., 2025) have significantly improved image-to-video (I2V) generation quality, they still struggle to faithfully preserve the identity of the input image, especially human faces. A significant challenge arises when the reference image contains a low-resolution human face and the motion prompt involves large-scale movements (e.g., jumping, turning). Under these conditions, as illustrated in Fig. 1, the propagation of errors across frames results in progressive identity degradation, causing the generated subject to drift from the appearance in the initial frame.
@@
We evaluate our methodology on both Wan 2.2 I2V model (Wan et al., 2025) and our in-house 15B I2V model using comprehensive quality and identity preservation metrics, comparing against state-of-the-art methods for identity-preserving video generation. Extensive experiments demonstrate that our approach generates high-quality videos with significantly improved identity consistency.
@@ 2 Related Work
Identity preservation in video generation. Current identity-preserving video generation methods mainly focus on T2V generation, relying on additional modules to inject identity information. ID-Animator (He et al., 2024) and ConsisID (Yuan et al., 2025) encode ID-relevant information with a face adapter. FantasyID (Zhang et al., 2025) introduces 3D facial geometry to preserve structure. Concat-ID (Zhong et al., 2025) and Stand-In (Xue et al., 2025a) concatenate reference face tokens at the end of video token and employ 3D self-attention to fuse them. Some methods (Wang et al., 2025; Xie et al., 2025; Song et al., 2025) propose mixture of facial experts to dynamically fuse identity, semantic, and detail via a adaptive gate. PersonalVideo (Li et al., 2024) utilizes per-scene LoRA adapters for video personalization, but it requires finetuning the pretrained model for each new person during inference. Unlike T2V generation, I2V models naturally support first-frame conditioning for better controllability. Adding extra modules to improve identity preservation would disrupt the original structure of foundation models and increase complexity. To address this, our method directly learns feedback from an identity reward function, producing videos that more faithfully meet identity-specific goals and maintain identity consistency.
@@ 5.1 Experimental setup
Implementation Details. We utilize both our in-house I2V model (15B dense model with both MMDiT block and single DiT block) ,Wan2.2 5B (Wan et al., 2025), and Wan2.2 27B-A14B (Wan et al., 2025) as our image-to-video foundational model. For Wan2.2 27B-A14B (Wan et al., 2025), we keep all modules frozen except the low-noise expert model. In order to improve training efficiency, we employ Wan2.2-Lightning, the distilled version of Wan2.2 (Wan et al., 2025) which requires only 8 steps without the need of Classifier-free guidance (CFG). For the training process, we employ the Adam optimizer configured with a learning rate of 2e-5, and train 100 steps with a batch size of 64. The truncation gradient step $K$ is $4$, the facial reward weight is $0.1$, and the KL-divergence loss weight is $1$.

### v2

@@ 1 Introduction
Identity preservation is a widely popular and explored topic in text-to-video (T2V) generation. Numerous recent methods (He et al., 2024; Liu et al., 2025c; Yuan et al., 2025; Wang et al., 2025b; Li et al., 2024; 2025; Song et al., 2025; Xie et al., 2025; Zhang et al., 2025; Zhong et al., 2025; Jiang et al., 2025) attempt to generate high-fidelity videos that maintain consistent identity from user-provided reference subjects, often by incorporating auxiliary identity modules and carefully curated datasets. However, it remains a persistent yet underexplored challenge in image-to-video (I2V) generation. Despite diffusion transformer-based models such as CogVideoX (Yang et al., 2024), HunyuanVideo (Kong et al., 2024), and Wan (Wan et al., 2025) have significantly improved image-to-video (I2V) generation quality, they still struggle to faithfully preserve the identity of the input image, especially human faces. A significant challenge arises when the reference image contains a low-resolution human face and the motion prompt involves large-scale movements (e.g., jumping, turning). Under these conditions, as illustrated in Fig. 1, the propagation of errors across frames results in progressive identity degradation, causing the generated subject to drift from the appearance in the initial frame.
@@
We evaluate our methodology on both Wan 2.2 I2V model (Wan et al., 2025) and our in-house 15B I2V model using comprehensive quality and identity preservation metrics, comparing against state-of-the-art methods for identity-preserving video generation. Extensive experiments demonstrate that our approach generates high-quality videos with significantly improved identity consistency. We leverage the powerful ROLL (Wang et al., 2025a) RL training framework in our implementationtiontiontiontion.
@@ 2 Related Work
Identity preservation in video generation. Current identity-preserving video generation methods mainly focus on T2V generation, relying on additional modules to inject identity information. ID-Animator (He et al., 2024) and ConsisID (Yuan et al., 2025) encode ID-relevant information with a face adapter. FantasyID (Zhang et al., 2025) introduces 3D facial geometry to preserve structure. Concat-ID (Zhong et al., 2025) and Stand-In (Xue et al., 2025a) concatenate reference face tokens at the end of video token and employ 3D self-attention to fuse them. Some methods (Wang et al., 2025b; Xie et al., 2025; Song et al., 2025) propose mixture of facial experts to dynamically fuse identity, semantic, and detail via a adaptive gate. PersonalVideo (Li et al., 2024) utilizes per-scene LoRA adapters for video personalization, but it requires finetuning the pretrained model for each new person during inference. Unlike T2V generation, I2V models naturally support first-frame conditioning for better controllability. Adding extra modules to improve identity preservation would disrupt the original structure of foundation models and increase complexity. To address this, our method directly learns feedback from an identity reward function, producing videos that more faithfully meet identity-specific goals and maintain identity consistency.
@@ 5.1 Experimental setup
Implementation Details. We utilize both our in-house I2V model (15B dense model with both MMDiT block and single DiT block), Wan2.2 5B (Wan et al., 2025) and Wan2.2 27B-A14B (Wan et al., 2025) as our image-to-video foundational model. For Wan2.2 27B-A14B (Wan et al., 2025), we keep all modules frozen except the low-noise expert model. We use ROLL (Wang et al., 2025a) as our training framework. In order to improve training efficiency, we employ Wan2.2-Lightning, the distilled version of Wan2.2 (Wan et al., 2025) which requires only 8 steps without the need of Classifier-free guidance (CFG). For the training process, we employ the Adam optimizer configured with a learning rate of 2e-5, and train 100 steps with a batch size of 64. The truncation gradient step $K$ is $4$, the facial reward weight is $0.1$, and the KL-divergence loss weight is $1$.